OpenStudy (anonymous):

Combine as indicated by the signs.4+a/a-3+4/5 i need help

OpenStudy (anonymous):

@Jonask

OpenStudy (anonymous):

@pinkmommy

OpenStudy (anonymous):

@richyw

OpenStudy (anonymous):

@shashawna

OpenStudy (anonymous):

4+a/a-3+4/5= 4a/7-a/5= 4a/-7a/5 -3a/5 can u figuit out from there?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

9a-8/-3a+5 i thik this answers

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

@Zarkon

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

@ninzinga

OpenStudy (anonymous):

@Fosterthepeople

OpenStudy (anonymous):

\[\frac{ a }{ b }+\frac{c}{d}=\frac{ad+bc}{bd}\]

OpenStudy (anonymous):

\[\frac{ 4+a }{ a-3 }+\frac{4}{5}=\frac{ 5(4+a)+4(a-3) }{ 5(a-3) }\]

OpenStudy (anonymous):

9a-8/-3a+5 i thik this answers

OpenStudy (anonymous):

is it the question like this fraction

OpenStudy (anonymous):

\[\frac{9a+8}{5a-15}\]